OpenStudy (ujjwal):

25 tuning forks are arranged in series in the order of decreasing frequency. Any two successive fork produce 3 beat/sec. If the frequency of first tuning fork is the octave of the last fork, the frequency of 21st fork is: a)72 Hz b)84 Hz c)87 Hz d)144 Hz

OpenStudy (vincent-lyon.fr):

84Hz?

OpenStudy (ujjwal):

No idea.. I don't have the answer.. well can you just say how do you get it?

OpenStudy (vincent-lyon.fr):

Difference between two successive frequencies is 3 Hz. Total is 24 intervals, so overall difference is 3 x 24 = 72 Hz. f25 doubled yields f1 and f25 + 72 Hz = f1 So f1 = 72 Hz and f25 = 144 Hz.

OpenStudy (ujjwal):

you mean f1=144 and f25=72..?? That makes f21=84.. Are you sure about the answer? I guess you are..

OpenStudy (vincent-lyon.fr):

I am sure that beat frequency between f1 and f2 is: | f1 - f2 | I am sure that an octave is f2/f1 = 2 It seems the rest follows from that, provided I have not made a mistake in counting the number of intervals!

OpenStudy (ujjwal):

yeah i guess you are right.. What i didn't know actually was the octave thing.. I didn't know how to relate first and last frequencies.. Thanks!

OpenStudy (vincent-lyon.fr):

@ujjwal For your information: Ascending octave is ratio 2:1 Asc. fifth is ratio 3:2 Asc. fourth is ratio 4:3 Asc. major third is ratio 5:4 Asc. minor third is ratio 6:5

OpenStudy (ujjwal):

those are the ratios of frequencies? @Vincent-Lyon.Fr

OpenStudy (vincent-lyon.fr):

Yep!

OpenStudy (anonymous):

\[\nu1-\nu21=60\] \[\nu1-\nu25=72\] are the given conditions \[\nu1=2*\nu25\] also the given data and using the above 3 equations \[so 2*\nu25-\nu25=72\] \[\nu25=72\] using this value in the second equation \[ \nu1=72+72\] =144